OpenStudy (anonymous):

Can anyone solve this using the method: Elimination.? 3x - 9y = 3 6x - 3y = -24 Thanks(:

OpenStudy (anonymous):

Don't you subtract everything by 2?

OpenStudy (anonymous):

yes -2*(3x - 9y = 3) -6x +18y=-6 6x - 3y = -24

OpenStudy (anonymous):

-6x +18y=-6 + 6x - 3y = -24 ---------------- 0 +15y=-30 y=-30/15= y=-2

OpenStudy (tyteen4a03):

Firstly find out the best way to have two variables match. We usually multiply an equation by a factor, but sometimes division will stop the number from getting larger, so choose wisely. (You cannot subtract, or add a value to everything: this is wrong!) Here, we'll multiply the first equation by 2 to make the second equation \(6x - 18y = 6\). Now, subtract the second equation from the first equation. You should get \(-15y = 30\). Solve for y and then use this newfound value of y to solve for x.

OpenStudy (anonymous):

6x - 3y = -24 => 6x - 3(-2) = -24 6x=-24+6 = -18 x=-18/6=-3

OpenStudy (anonymous):

so x=-3 y=-2

OpenStudy (anonymous):

How you solve for x and y.?

OpenStudy (tyteen4a03):

Try reading my explanation and @baziham 's actual work.

OpenStudy (anonymous):

multip -2 in eqution 1 then sum two equations with them you get 0 +15y=-30 and by replacement "y =-2" in equation 1 (or two) you get x value

OpenStudy (anonymous):

Oh ok, Thank you guys so much! You're such a big helpp.(:

OpenStudy (anonymous):

your welcome